# STCast: Adaptive Boundary Alignment for Global and Regional Weather Forecasting

### v1

@@ Abstract
To gain finer regional forecasts, many works have explored the regional integration from the global atmosphere, e.g., by solving boundary equations in physics-based methods or cropping regions from global forecasts in data-driven methods. However, the effectiveness of these methods is often constrained by static and imprecise regional boundaries, resulting in poor generalization ability. To address this issue, we propose Spatial-Temporal Weather Forecasting (STCast), a novel AI-driven framework for adaptive regional boundary optimization and dynamic monthly forecast allocation. Specifically, our approach employs a Spatial-Aligned Attention (SAA) mechanism, which aligns global and regional spatial distributions to initialize boundaries and adaptively refines them based on attention-derived alignment patterns. Furthermore, we design a Temporal Mixture-of-Experts (TMoE) module, where atmospheric variables from distinct months are dynamically routed to specialized experts using a discrete Gaussian distribution, enhancing the model’s ability to capture temporal patterns. Beyond global and regional forecasting, we evaluate our STCast on extreme event prediction and ensemble forecasting. Experimental results demonstrate consistent superiority over state-of-the-art methods across all four tasks.
Introduction
@@
Why need global forecasts to support regional forecasting? Achieving accurate, kilometre-scale regional weather forecasting is still a formidable scientific task with far-reaching socio-economic impact. Existing strategies typically fall into two paths: training a dedicated regional model or extracting the regional prediction from a global forecast. Traditional Numerical Weather Prediction (NWP) (Bauer, Thorpe, and Brunet 2015; Lynch 2008; Kalnay 2002) methods solve partial differential equations (PDEs) at finer resolutions, but incur prohibitively high computational cost. Recent data-driven approaches (Bi et al. 2023; Chen et al. 2023b; Lam et al. 2023; Bodnar et al. 2025; Subich et al. 2025; Wu et al. 2025) significantly reduce this cost by neural networks. However, these models often rely on patch embeddings that downsample input variables, resulting in the loss of fine-grained local details. That is, training a global model at 1 km resolution (approximately 0.01°, or ) would be computationally infeasible. Conversely, restricting training to a high-resolution region neglects cross-regional dependencies that are critical for accurate forecasting. Thus, both direct training of high-resolution regional models and extracting them from high-resolution global forecasts are impractical. These limitations highlight the need for hybrid frameworks that couple low-resolution global forecasts with high-resolution regional forecasting in a computationally efficient manner.
@@
Why adaptive boundaries must cover the global area? While hybrid global–regional frameworks are gaining attention, existing coupling strategies, whether based on traditional NWP (Lundquist, Chow, and Lundquist 2010; Mani 2012) or AI models (Gao et al. 2025; Adamov et al. 2025; Nipen et al. 2024), typically define regional boundaries using only adjacent areas. This local perspective contradicts the well-established Atmosphere–Ocean–Land–Biosphere Coupling Theory (Manabe and Bryan 1969; Zhang, Tian, and Wang 2018), which posits that any points in the regional atmosphere are influenced by the entire Earth system. For example, Siberian cold surges can trigger East-Asian cold waves, and surface heating over the Tibetan Plateau can simultaneously alter East Asian monsoons and North American jet stream (Wu et al. 2023). Thus, the true boundary for a region is not its neighbors, but the entire Earth.
@@
Spatial-Aligned Attention (SAA) incorporates a learnable global-regional distribution into linear cross-attention, enabling adaptive aggregation of global atmospheric information for regional forecasting. To couple the global and regional variables, SAA employs two key mechanisms: (1) a Manhattan distance metric to quantify spatial separation from the target region, and (2) an exponential distance-decay function to initialize the learnable global-regional distribution, ensuring weaker influence from distant regions. This prior modulates the attention weights by element-wise multiplication and is further refined during training. As a result, the global–regional correlation evolves dynamically, aligning spatial dependencies with physical intuition throughout the optimization process.
@@
Related work
Global-Regional Weather Coupling
@@
Accurate global–regional coupling remains a core challenge in regional forecasting due to the difficulty of boundary specification. NWP models address this by solving PDEs under prescribed boundary, e.g., sponge layers (Mani 2012), Dirichlet (Hidayatullah et al. 2019), and Neumann conditions (Sabathier et al. 2023). In contrast, AI approaches (Gao et al. 2025; Adamov et al. 2025; Nipen et al. 2024) bypass boundary equations by concatenating a fixed neighborhood in global area to region, resulting in static and local coupling.
@@
The recent work, OneForecast (Gao et al. 2025), tackles the same 4 tasks as ours by concatenating neighboring low-resolution global forecasts with high-resolution regional variables. In contrast, our method replaces static concatenation with a transformer-based framework that adaptively models global–regional correlations, guided by a learned prior. This enables dynamic boundary refinement and long-range dependency modeling beyond local neighborhoods.
Data-Driven Weather Forecasting
@@
Prior to deep learning, numerical weather prediction (NWP) prevailed, producing forecasts by solving PDEs on high-resolution global grids (Bauer, Thorpe, and Brunet 2015; Lynch 2008; Kalnay 2002). These methods deliver physically analysis and rigorously validated forecasts (Molteni et al. 1996; Ritchie et al. 1995), but their computation at inference remains prohibitively high, especially at fine scales.
@@
The emergence of large-scale atmospheric reanalyses has catalysed a shift toward data-driven forecasting. Early exemplars, FourCastNet (Kurth et al. 2023) and Pangu-Weather (Bi et al. 2023), employ Fourier Neural Operators (FNO) (Li et al. 2021) and 3D Swin Transformers (Liu et al. 2021), respectively, to approximate atmospheric evolution. Subsequent research has bifurcated into two streams: (i) neural operators, such as KNO (Xiong et al. 2023) and SFNO (Bonev et al. 2023), that directly learn the temporal evolution operator; and (ii) neural networks, including FengWu (Chen et al. 2023a), FengWu-ghr (Han et al. 2024), GraphCast (Lam et al. 2023), FuXi (Chen et al. 2023b), GenCast (Price et al. 2025), and Stormer (Nguyen et al. 2025), which leverage inductive biases tailored to atmosphere. All achieve competitive skill with limited computation.
@@
Inspired by the success of MoE in LLMs (Shazeer et al. 2017), recent studies have begun to integrate MoE into weather forecasting. VAMoE (Chen et al. 2025) extends MoE to incremental weather forecasting; EWMoE (Gan et al. 2025) augments FourCastNet with MoE layers. In contrast, we propose TMoE that explicitly partitions atmospheric inputs by month and dynamically routes them to specialized temporal experts. This structure enables STMoE to capture inter-month variability and intra-month correlation.
Methodology
@@
We give a unified framework for 4 tasks: low-resolution global prediction, high-resolution regional prediction, typhoon track forecasting, and ensemble forecasting. Section Overview formally defines each task. We then introduce Spatial-Aligned Attention (SAA) that fuses global and regional atmospheric variables via a learnable global-regional distribution, and Temporal Mixture-of-Experts (TMoE) that allocates monthly data to specialized experts.
Overview
@@
For the weather forecasting task, the AI model infers future atmospheric states from historical fields , i.e., . Here, comprises upper-air variables on 13 pressure levels and surface variables , where and denote the number of variables per pressure and surface level, respectively.
@@
As illustrated in Fig. 2, a framework unifies four subtasks: global deterministic forecasting , high-resolution regional forecasting , typhoon track prediction , and ensemble forecasting . In Fig. 2(a), we introduce Temporal Mixture-of-Experts (TMoE) and further integrate Flash-Attention (Dao et al. 2022) with MoE (Shazeer et al. 2017). Recognizing the pronounced seasonal variability of atmospheric states, TMoE treats monthly forecast as a distinct tasks and assigns it to multiple dedicated experts. Global forecasting is expressed as , where denotes the global variables. For regional forecasting, we employ a global–regional coupling strategy in Fig. 2(b). A Spatial-Aligned Attention (SAA) module fuses global variables with regional inputs to yield high-resolution predictions: . The predicted MSL is subsequently used to infer typhoon tracks in Fig. 2(c). Beyond deterministic forecasts, we evaluate probabilistic skill for both long-range and ensemble scenarios in Fig. 2(d). Perlin noise is injected into initial state , and the model is run times; the ensemble mean provides probabilistic forecast.
@@
Spatial-Aligned Attention
@@
For precise quantification of spatial relationships, SAA calculates Manhattan distance between each global point and target region. This distance metric is defined as:
where denotes the center coordinates of target region, and represent its height and width. This efficient formulation introduces negligible overhead.
@@
Model Metric 6-hour 1-day 4-day 7-day 10-day RMSE ACC RMSE ACC RMSE ACC RMSE ACC RMSE ACC Pangu-weather(Bi et al. 2023) 0.0826 0.9876 0.1571 0.9581 0.3380 0.8167 0.5092 0.5738 0.6215 0.3542 Graphcast(Lam et al. 2023) 0.0626 0.9928 0.1304 0.9705 0.2861 0.8705 0.4597 0.6692 0.6009 0.4275 Fuxi(Chen et al. 2023b) 0.0987 0.9820 0.1708 0.9511 0.4128 0.7379 0.5972 0.4446 0.6981 0.2391 Oneforecast(Gao et al. 2025) 0.0549 0.9943 0.1231 0.9737 0.2732 0.8825 0.4468 0.6888 0.5918 0.4457 Ours 0.0617 0.9956 0.1197 0.9740 0.2578 0.8927 0.4348 0.7019 0.5763 0.4715
Temporal Mixture-of-Experts
@@
Ablation Studies Metric 6-hour 1-day 4-day 7-day 10-day RMSE ACC RMSE ACC RMSE ACC RMSE ACC RMSE ACC High-resolution Regional Forecasts w/o SAA 0.0767 0.9762 0.1802 0.8675 0.4001 0.4127 0.5297 0.4229 0.7610 0.2541 w/o Global-Regional Distribution 0.0694 0.9805 0.1631 0.8864 0.3794 0.6718 0.5082 0.4566 0.7192 0.3286 w SAA 0.0493 0.9946 0.0854 0.9854 0.2068 0.9203 0.3712 0.7442 0.4945 0.5433 Low-resolution Global Forecasts w/o TMoE 0.0751 0.9915 0.1451 0.9714 0.3249 0.8201 0.5109 0.5412 0.6426 0.3184 w/o Month Embedding 0.0744 0.9928 0.1346 0.9764 0.2865 0.8180 0.4631 0.5941 0.6049 0.3559 w TMoE 0.0617 0.9956 0.1197 0.9740 0.2578 0.8927 0.4348 0.7019 0.5763 0.4715
2024.11.04, T+0
2024.11.06, T+12
2024.11.06, T+12
@@
Experiments
Main Results
Low-resolution Global Weather Forecasting. We evaluate forecasting performance using two standard metrics: RMSE and ACC. Due to significant scale variations across atmospheric variables, direct comparison using absolute values is infeasible. We therefore present normalized RMSE and ACC scores in Tab. 1, where STCast demonstrates consistent superiority over baselines across all benchmarks, with particularly significant gains in long-term predictions. Further validation through real-value RMSE and ACC comparisons (1-10 day) in Fig. 5 confirms STCast’s state-of-the-art performance across multiple variables. This enhancement is attributed to our month-specific training strategy, which effectively captures both seasonal dependencies and month-to-month variations in weather systems. Complementary visualization in Fig. 6 compares spatial error distributions for five key variables across three methods, providing qualitative evidence of STCast’s reduced prediction uncertainty.
High-resolution Regional Weather Forecasting. As demonstrated in Fig. 1, we compare mean RMSE and ACC scores across five surface variables against Direct Training and OneForecast (Gao et al. 2025). Quantitative analysis reveals that direct-trained STCast (without dynamic boundary) and OneForecast achieve comparable performance. However, implementing our dynamic boundary condition in STCast yields significant improvements: mean RMSE decreases by 0.05 while mean ACC increases by 0.1. This enhancement confirms the critical role of adaptive boundary modeling in regional forecasting systems. Complementary visualization of 6-hour regional forecasts for U10 and MSL in Fig. 8 provides performance validation. Error analysis demonstrates STCast’s superior accuracy, achieving near-zero relative errors of just 0.7% for U10 and 0.1% for MSL - substantially lower than competitors.

### v2

@@ Abstract
To gain finer regional forecasts, many works have explored the regional integration from the global atmosphere, e.g., by solving boundary equations in physics-based methods or cropping regions from global forecasts in data-driven methods. However, the effectiveness of these methods is often constrained by static and imprecise regional boundaries, resulting in poor generalization ability. To address this issue, we propose Spatial-Temporal Weather Forecasting (STCast), a novel AI-driven framework for adaptive regional boundary optimization and dynamic monthly forecast allocation. Specifically, our approach employs a Spatial-Aligned Attention (SAA) mechanism, which aligns global and regional spatial distributions to initialize boundaries and adaptively refines them based on attention-derived alignment patterns. Furthermore, we design a Temporal Mixture-of-Experts (TMoE) module, where atmospheric variables from distinct months are dynamically routed to specialized experts using a discrete Gaussian distribution, enhancing the model’s ability to capture temporal patterns. Beyond global and regional forecasting, STCast is evaluated on extreme event prediction and ensemble forecasting. Experimental results demonstrate consistent superiority over state-of-the-art methods across all four tasks.
1 Introduction
@@
Why do we need global forecasts to support regional forecasting? Achieving accurate, kilometre-scale regional weather forecasting is still a formidable scientific task with far-reaching socio-economic impact. Existing strategies typically fall into two paths: training a dedicated regional model or extracting the regional prediction from a global forecast. Traditional Numerical Weather Prediction (NWP) [2, 27, 19] methods solve partial differential equations (PDEs) at finer resolutions, but incur prohibitively high computational cost. Recent data-driven approaches [3, 8, 22, 4, 41, 45] significantly reduce this cost by neural networks. However, these models often rely on patch embeddings that downsample input variables, resulting in the loss of fine-grained local details. That is, training a global model at 1 km resolution (approximately 0.01°, or ) would be computationally infeasible. Conversely, restricting training to a high-resolution region neglects cross-regional dependencies that are critical for accurate forecasting. Thus, both direct training of high-resolution regional models and extracting them from high-resolution global forecasts are impractical. These limitations highlight the need for hybrid frameworks that couple low-resolution global forecasts with high-resolution regional forecasting in a computationally efficient manner.
@@
Why must adaptive boundaries cover the global area? While hybrid global–regional frameworks are gaining attention, existing coupling strategies, whether based on traditional NWP [26, 30] or AI models [11, 1, 34], typically define regional boundaries using only adjacent areas. This local perspective contradicts the well-established Atmosphere–Ocean–Land–Biosphere Coupling Theory [29, 52], which posits that any points in the regional atmosphere are influenced by the entire Earth system. For example, Siberian cold surges can trigger East-Asian cold waves, and surface heating over the Tibetan Plateau can simultaneously alter East Asian monsoons and North American jet stream [46]. Thus, the true boundary for a region is not its neighbors, but the entire Earth.
@@
Spatial-Aligned Attention (SAA) incorporates a learnable global-regional distribution into linear cross-attention, enabling adaptive aggregation of global atmospheric information for regional forecasting. To couple the global and regional variables, SAA employs two key mechanisms: (1) a Great Circle distance metric to quantify spatial separation from the target region, and (2) an exponential distance-decay function to initialize the learnable global-regional distribution, ensuring weaker influence from distant regions. This prior modulates the attention weights by element-wise multiplication and is further refined during training. As a result, the global–regional correlation evolves dynamically, aligning spatial dependencies with physical intuition throughout the optimization process.
@@
2 Related work
2.1 Global-Regional Weather Coupling
@@
Accurate global–regional coupling remains a core challenge in regional forecasting due to the difficulty of boundary specification. NWP models address this by solving PDEs under prescribed boundary, e.g., sponge layers [30], Dirichlet [17], and Neumann conditions [37]. In contrast, AI approaches [11, 38, 1, 34] bypass boundary equations by concatenating a fixed neighborhood in global area to region, resulting in static and local coupling.
@@
The recent work, OneForecast [11], tackles the same 4 tasks as ours by concatenating neighboring low-resolution global forecasts with high-resolution regional variables. In contrast, our method replaces static concatenation with a transformer-based framework that adaptively models global–regional correlations, guided by a learned prior. This enables dynamic boundary refinement and long-range dependency modeling beyond local neighborhoods.
2.2 Data-Driven Weather Forecasting
@@
Prior to deep learning, NWP prevailed, producing forecasts by solving PDEs on high-resolution global grids [2, 27, 19]. These methods deliver physically analysis and rigorously validated forecasts [31, 36], but their computation at inference remains prohibitively high, especially at fine scales.
@@
The emergence of large-scale atmospheric reanalyses has catalysed a shift toward data-driven forecasting. Early exemplars, FourCastNet [21] and Pangu-Weather [3], employ Fourier Neural Operators (FNO) [23] and 3D Swin Transformers [24], respectively, to approximate atmospheric evolution. Subsequent research has bifurcated into two streams: (i) neural operators, such as KNO [49] and SFNO [5], that directly learn the temporal evolution operator; and (ii) neural networks, including FengWu [7], FengWu-ghr [14], Graphcast [22], FuXi [8], GenCast [35], and Stormer [33], which leverage inductive biases tailored to atmosphere. All achieve competitive performance.
@@
Inspired by the success of MoE in LLMs [39], recent studies have begun to integrate MoE into weather forecasting. VAMoE [6] extends MoE to incremental weather forecasting; EWMoE [10] augments FourCastNet with MoE layers. In contrast, we propose TMoE that explicitly partitions inputs by month and dynamically routes them to specialized temporal experts. This structure enables TMoE to capture inter-month variability and intra-month correlation.
3 Methodology
@@
We give a unified framework for 4 tasks: low-resolution global prediction, high-resolution regional prediction, typhoon track forecasting, and ensemble forecasting. Sec. 3.1 formally defines each task. We then introduce Sec. 3.2 that fuses global and regional atmospheric variables via a learnable global-regional distribution, and Sec. 3.3 that allocates monthly data to specialized experts.
3.1 Overview
STCast is composed of three components: the Encoder, Processor, and Decoder. The Processor employs an alternating strategy that integrates window-based attention with self-attention. This hybrid design allows the model to capture both local and global dependencies within the input distribution. More details are provided in the Appendix.
@@ 3.1 Overview
This framework is applied to the weather forecasting task, where the model predicts future atmospheric states based on historical inputs . Specifically, , where includes upper-air variables across 13 pressure levels and surface variables , with N and M denoting the number of variables per pressure and surface level, respectively.
@@
As illustrated in Fig. 2, STCast unifies four key subtasks to address diverse forecasting challenges: global deterministic forecasting , high-resolution regional forecasting , typhoon track prediction , and ensemble forecasting . To enhance temporal modeling, we introduce the Temporal Mixture-of-Experts (TMoE) and integrate Flash-Attention [9] with MoE [39], as shown in Fig. 2(a). Considering the significant seasonal variability in atmospheric dynamics, TMoE assigns monthly forecasts to specialized experts. For global forecasting, this strategy is expressed as , where denotes the global input variables. In regional forecasting, we adopt a global–regional coupling approach as depicted in Fig. 2(b). The Spatial-Aligned Attention (SAA) module fuses global and regional inputs to produce high-resolution predictions: . The predicted mean sea level pressure (MSL) is subsequently used to infer typhoon tracks illustrated in Fig. 2(c). Beyond deterministic tasks, we also evaluate probabilistic forecasting for long-range and ensemble scenarios as shown in Fig. 2(d). To this end, we inject Perlin noise into the initial global state and execute ensemble simulation for n times. The ensemble mean is used to provide probabilistic predictions.
@@
3.2 Spatial-Aligned Attention
@@
To achieve precise quantification of spatial relationships, SAA calculates Great Circle distance between each global point and target region. Compared to Manhattan and Euclidean distance, the Great Circle distance provides a more accurate measure of spatial relationships on the Earth’s surface. This distance metric is defined as:
where represent the latitude and longitude coordinates, denote the latitude and longitude of the target region’s center, define the region’s height and width, and indicate the distances in the latitude and longitude directions, respectively. This efficient formulation introduces negligible overhead.
@@
Regarding the cross-attention mechanism, SAA incorporates an efficient attention [40] to compute global-regional correlations. This design reduces computational complexity from to . The process is formulated as:
where denotes the embedding dimension of Key.
@@
3.3 Temporal Mixture-of-Experts
@@
4 Experiments
Model 6-hour 1-day 4-day 7-day 10-day RMSE ACC RMSE ACC RMSE ACC RMSE ACC RMSE ACC Pangu-weather[3] 0.0826 0.9876 0.1571 0.9581 0.3380 0.8167 0.5092 0.5738 0.6215 0.3542 Graphcast[22] 0.0626 0.9928 0.1304 0.9705 0.2861 0.8705 0.4597 0.6692 0.6009 0.4275 Fuxi[8] 0.0987 0.9820 0.1708 0.9511 0.4128 0.7379 0.5972 0.4446 0.6981 0.2391 OneForecast[11] 0.0549 0.9943 0.1231 0.9737 0.2732 0.8825 0.4468 0.6888 0.5918 0.4457 Ours 0.0617 0.9956 0.1197 0.9740 0.2578 0.8927 0.4348 0.7019 0.5763 0.4715
4.1 Dataset and Implementation Details
In this work, we conduct experiments on a popular weather dataset, i.e., ERA5 [16], provided by the ECMWF [31]. ERA5 dataset is a reanalysis atmospheric dataset, consisting of the atmospheric variables from 1979 to the present day with a 0.25°spatial resolution with . The global model is trained on 40-year dataset from 1979 to 2019 with 70 variables, and the regional model is trained in the same period with 5 surface variables on Eastern Asia.
We apply the AdamW optimizer with 0.0002 learning rate to model training. In both global and regional forecasts, we train 100 epochs and set batch size to 16. Our model is trained with 16 NVIDIA Tesla A100 GPUs.
Statement. More details about Dataset and Implementation Details, Numerical Comparison and Computation Metric of typhoon track prediction, computation comparisons, additional results, additional visualization, and additional ablation studies are provided in Appendix.
Ablation Studies 6-hour 1-day 4-day 7-day 10-day RMSE ACC RMSE ACC RMSE ACC RMSE ACC RMSE ACC High-resolution Regional Forecasts w/o SAA 0.0767 0.9762 0.1802 0.8675 0.4001 0.4127 0.5297 0.4229 0.7610 0.2541 w/o Global-Regional Distribution 0.0694 0.9805 0.1631 0.8864 0.3794 0.6718 0.5082 0.4566 0.7192 0.3286 w SAA 0.0478 0.9946 0.0845 0.9854 0.2051 0.9203 0.3699 0.7442 0.4921 0.5433 Low-resolution Global Forecasts w/o TMoE 0.0751 0.9915 0.1451 0.9714 0.3249 0.8201 0.5109 0.5412 0.6426 0.3184 w/o Month Embedding 0.0744 0.9928 0.1346 0.9764 0.2865 0.8180 0.4631 0.5941 0.6049 0.3559 w TMoE 0.0617 0.9956 0.1197 0.9740 0.2578 0.8927 0.4348 0.7019 0.5763 0.4715
@@ 4.1 Dataset and Implementation Details
4.2 Main Results
Low-resolution Global Forecasts. We evaluate the performance with two metrics: RMSE and ACC. Due to significant scale variations across atmospheric variables, direct comparison using absolute values is infeasible. We therefore present normalized RMSE and ACC scores in Tab. 1, where STCast demonstrates consistent superiority over baselines across all benchmarks, with particularly significant gains in long-term predictions. Further validation through real-value RMSE and ACC comparisons (1-10 day) in Fig. 5 confirms STCast’s state-of-the-art performance across multiple variables. This enhancement is attributed to our month-specific training strategy, which effectively captures both seasonal dependencies and month-to-month variations in atmospheric system. Complementary visualization in Fig. 6 compares spatial error distributions for five key variables across three methods, providing qualitative evidence of STCast’s reduced prediction uncertainty.
High-resolution Regional Forecasts. As demonstrated in Fig. 1, we compare mean RMSE and ACC scores across five surface variables against Direct Training and OneForecast [11]. Quantitative analysis reveals that direct-trained STCast (without dynamic boundary) and OneForecast achieve comparable performance. However, implementing our dynamic boundary condition in STCast yields significant improvements: RMSE decreases by 0.05 while ACC increases by 0.1. This enhancement confirms the critical role of adaptive boundary modeling in regional forecasting systems. Complementary visualization of 6-hour regional forecasts for X-direction wind at 10m height (U10) and Mean Sea-Level pressure (MSL) in Fig. 8 provides validation. Error analysis demonstrates STCast’s superior accuracy, achieving near-zero errors of 0.7% for U10 and 0.1% for MSL - substantially lower than competitors.
Extreme Events Assessment. Extreme weather events, particularly tropical cyclones, pose significant societal risks that demand accurate forecasting capabilities [43, 44]. To evaluate STCast’s performance under such critical conditions, we analyze two recent typhoon events: Typhoon Ewiniar (May 2024) and Typhoon Yinxing (November 2024) [51, 25]. As visualized in Fig. 9a-b, STCast’s 72-hour track forecasts demonstrate substantially closer alignment with observed paths compared to ECMWF [31], FourCastNet [21], Pangu-Weather [3], and FengWu [7] for both systems. Quantitative analysis in Fig. 9c reveals that while STCast achieves comparable performance to others in short-term forecasts, it significantly outperforms them in long-term predictions for Typhoon Yinxing, attaining a mean error of only 96.5 km compared to 160 km for the next-best Pangu-Weather. This improved accuracy, coupled with STCast’s consistent capability to capture cyclones’ structural development and translational dynamics, highlights its strong potential for extreme event forecasting.